OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2

OpenStudy (texaschic101):

2x + y = 1 -- subtract 2x from both sides y = -2x + 1 now sub -2x + 1 in for y in the other equation to find x. Give it a try.

OpenStudy (texaschic101):

let me start you off... 4x + 2(-2x + 1) = -2

OpenStudy (anonymous):

-8x + 3= -2?

OpenStudy (anonymous):

@texaschic101